OpenStudy (anonymous):

Find the equation of the line passing through the points (-12,-24) and (15,48). Write your answer in the y=mx+b form

OpenStudy (anonymous):

find the salope \[\Large m=\frac{y_{2}-y_{1}}{x_{2}-x_{1}}=\frac{48-(-24)}{15-(-12)}=\frac{72}{27}\] then you point slope form \[\Large y-y_{1}=m(x-x_{1})\] where m=72/27 y1=48 x1=15 use these values in the above equation to get the required equation of lines